40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2495     Entry time: Fri May 1 13:27:07 2020
Author: Raymond  
Type: Update 
Category: PD noise 
Subject: Sb3513 A2P6 2020-02-04 Dark Noise/QE data 

Attachment 1 contains the SR785 dark noise measurements at number of PD reverse bias voltages from 77-295K with filenaming convention:

[PD]_drkspec_[date]_[temp]_[input V]_[scan freq]_[FEMTO gain]_[date]_.txt

It also contains the keithley sweeps for QE calculations.

I'm still working out what is wrong with the QE data and how to effectively process the dark noise versus temperature. 

Attachment 1: 20200204_A2P6_77Kto295K.zip  768 kB  Uploaded Fri May 1 14:28:10 2020
ELOG V3.1.3-